OpenStudy (anonymous):

someone pls help ill medal but help ive been stuck forever What is the slope, and what is the y-intercept of the line? y=5/4x=9 someone pls help

OpenStudy (kiakat):

the y intercept is where is crosses the y axis, correct?

OpenStudy (kiakat):

y=mx+b ^ is the slope y=mx+b ^ is the y-intercept

OpenStudy (anonymous):

5/4 is slope and 9 is y-intercept

OpenStudy (anonymous):

The reason being is because y=mx+b m=slope in this case it is 5/4 b= y-intercept in this case 9 so y= 5/4x+9 m(slope)=5/4 and b(y-intercept)=9 So if you need it again you can do it without any problem! :D

OpenStudy (anonymous):

@igivemedalsandfollow! are you getting it?